9. Two boys move a crate 3m long and exert on it the forces F1 = F2 = 300N, forming a pair of forces which form an angle of 30⁰ if the force would be in the vertical direction straight. Find out what is the moment of the couple of forces that the two boys exercise at the box office.


1
Expert's answer
2021-03-29T09:04:38-0400
"M=2Fd\sin{30}\\M=2(300)(3)\sin{30}=900\ N"